You observe that demented ice cream truck phenomenon one day. When the truck is stationary nearby, it is playing a song of
Answer #1

by doppler effect

f' = f * ( V±V0 ) /(V±Vs)

f' = frequency observed by observer

f= frequency emitted by source

V = speed of sound = 343 m/s

V0 = observer speed and Vs   source speed

so  

f' = f * ( V±V0 ) /(V±Vs)

543 = 550* ( 343 +0 ) /( 343 +Vs)

543 = 550 343 + 0 343 + V

543 = - 188650 V + 343

543 v = 2401 for # -343)

from here Vs =4.4217 m/s answer

****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..